if f(z) is an analytical function of z, then prove that ( d^2/dx^2Β + d^2/ dy^2) log |f(z)|= 0


1
Expert's answer
2022-01-31T16:32:17-0500

We know that "\log |f| = \frac{1}{2} \log(f\bar{f})". If we write "f=u+iv", where "u, v" are real functions, we find "\log(f\bar{f})=\log (u^2+v^2)".

Therefore, we have

"(\frac{\partial^2}{\partial x^2}+\frac{\partial^2}{\partial y^2}) \log|f| = \frac{1}{2} (\frac{\partial^2}{\partial x^2}+\frac{\partial^2}{\partial y^2})\log(u^2+v^2)" and let us calculate the derivatives :

"\frac{1}{2}\frac{\partial^2}{\partial x^2} \log(u^2+v^2)=\partial_x(\frac{u\partial_x u + v\partial_x v}{u^2+v^2})" taking the second derivative gives us a long expression :

"(u^2+v^2)(u\partial_{xx}u +v\partial_{xx} v+(\partial_xu)^2+(\partial_xv)^2)-\\ -2u^2(\partial_xu)^2-2v^2(\partial_xv)^2-4uv\partial_xu\partial_xv" all divided by "(u^2+v^2)^2"

Developping this expression gives us

"\frac{(u^2+v^2)(u\partial_{xx}u+v\partial_{xx} v) + (v\partial_x u - u\partial_x v)^2-(u\partial_x u + v\partial_x v)^2}{(u^2+v^2)^2}"

Calculating the second derivative with respect to "y" gives us a similar expression :

"\frac{(u^2+v^2)(u\partial_{yy}u+v\partial_{yy} v) + (v\partial_y u - u\partial_y v)^2-(u\partial_y u + v\partial_y v)^2}{(u^2+v^2)^2}"

Now let us use the Cauchy-Riemann equations (as "f" is analytic, "u, v" should satisfy them) :

"\begin{cases} \partial_x u = \partial_y v \\ \partial_y u = -\partial_x v \end{cases}" and therefore "(\frac{\partial^2}{\partial x^2}+\frac{\partial^2}{\partial y^2})u = (\frac{\partial^2}{\partial x^2}+\frac{\partial^2}{\partial y^2})v=0"

In addition, we have "(u\partial_y u + v\partial_y v)^2 = (v\partial_x u - u\partial_x v)^2" and "(u\partial_x u+v\partial_x v)^2 = (u\partial_y v - v\partial_y u)^2" by Cauchy-Riemann equations. Therefore, the sum

"(\frac{\partial^2}{\partial x^2}+\frac{\partial^2}{\partial y^2}) \log|f| = \frac{1}{2} (\frac{\partial^2}{\partial x^2}+\frac{\partial^2}{\partial y^2})\log(u^2+v^2)=0"